40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 13011     Entry time: Wed May 24 18:19:15 2017
Author: Kaustubh 
Type: Update 
Category: General 
Subject: ET-3010 PD Test 

Summary:

In continuation to the previous test conducted on the ET-3040 PD,  I performed a similar test on the ET-3010 model. This model requires a fiber couple input for proper testing, but I tested it in free space without a fiber couple as the laser power was only 1.00 mW and there was not much danger of scattering of the laser beam. The Data Sheet can be found here

Procedure:

The schematic(attached below) and the procedure are the same as the previous time. The pump current was set to 19.5 mA giving us a laser beam of power 1.00mW at the fiber couple output. The measured voltage for the reference detector was 1.8V. For the DUT, the voltage is amplified using a low noise amplifier(model SR-560) with a gain of 100. Without any laser incidence on the DUT, the multimeter reads 120.6 mV. After alligning the laser with the DUT, the multimeter reads 348.5 mV, i.e. the voltage for the DUT is 227.9/100 ~ 2.28 mV. The DC transimpedance of the reference detector is 10kOhm and its responsivity to 1064 nm is around 0.75 A/W. Using this we calculate the power at the reference detector to be 0.24 mW. The DC transimpedance for the DUT is 50Ohm and the responsivity is around 0.85 A/W. Using this we calculate the power at the DUT to be 0.054 mW. After this we connect the the laser input to the Netwrok Analyzer(AG4395A) and give an RF signal with -10dBm and frequency modulation from 100 kHz to 500 MHz.The RF output from the Analyzer is coupled to the Reference Channel(CHR) of the analyzer via a 20dB directional coupler. The AC output of the reference detector is given at Channel A(CHA) and the output from the DUT is given to Channel B(CHB). We got plots of the ratios between the reference detector, DUT and the coupled refernce for the Transfer Function and the Phase. I stored the data under the directory.../scripts/general/netgpibdata/data. The Bode Plot has been attached below and seeing it we observe that the cut-off frequency for the ET-3010 model is atleast over 500 MHz(stated as >1.5 GHz in the data sheet).

Result:

The bandwidth of the ET-3010 PD is atleast 500MHz, stated in the data sheet as >1.5GHz.

Precaution:

The ET-3010 PD has an internal power supply of 6V. Don't leave the PD connected to any instrument after the experimentation is done or else the batteries will get drained if there is any photocurrent on the PDs.

To Do:

Caliberate the vertical axis in the Bode Plot with transimpedance(Ohms) for the two PDs. Automate the procedure by making a Python script for taking multiple set of readings from the Netwrok Analyzer and aslo plot the error bands.

Attachment 1: PD_test_setup.png  41 kB  | Hide | Hide all
PD_test_setup.png
Attachment 2: ET-3010_test.pdf  38 kB  Uploaded Thu May 25 00:30:49 2017  | Hide | Hide all
ET-3010_test.pdf
Attachment 3: ET-3010_test.zip  12 kB  Uploaded Sat May 27 13:11:41 2017
ELOG V3.1.3-